# MAC-VO: Metrics-aware Covariance for Learning-based Stereo Visual Odometry mac-vo.github.io

### v1

@@ IV-A Quantitative Analysis
TABLE I: Performance comparison of different methods on the EuRoC Dataset. Only odd-ordered trajectory is shown due to page limit, see Appendix. B-A for the remaining results.
@@
      ORB-SLAM 3 0.0035 0.0450 0.0058 0.0603 0.0059 0.0526 0.0096 0.1757 0.0064 0.1615 0.033 0.9497 0.0092 0.1866
      DROID-SLAM 0.0012 0.0159 0.0034 0.2656 0.0025 0.0193 0.0026 0.0417 0.0012 0.0289 0.0034 0.1033 0.0024 0.0590
@@
      TartanVO⋆† 0.0121 0.0560 0.0302 0.2791 0.0193 0.0604 0.0251 0.1244 0.0065 0.0920 0.0303 0.2986 0.0198 0.1270
      TartanVO 0.0277 0.5122 0.0514 0.6635 0.0464 0.4797 0.0394 1.0420 0.0195 0.4684 0.0473 1.9657 0.0368 0.8346
      iSLAM-VO 0.0042 0.0560 0.0076 0.2789 0.0070 0.0603 0.0066 0.1241 0.004 0.0920 0.0151 0.2984 0.0071 0.1269
      DPVO⋆† 0.0015 0.0207 0.0028 0.0273 0.0028 0.0243 0.0041 0.0496 0.0016 0.0342 0.0045 0.1205 0.0027 0.0437
Ours 0.0014 0.0214 0.0023 0.0238 0.0025 0.0216 0.0029 0.0434 0.0012 0.0289 0.0049 0.1284 0.0024 0.0403
Average over all trajectories of EuRoC.   Monocular method.   Scale-aligned with ground truth.
EuRoC Dataset    We assessed our model on the EuRoC [45] dataset, as detailed in Table. I, comparing it against baseline methods including visual odometries and state-of-the-art visual SLAM systems with loop-closure and global bundle adjustment. While our method exhibits compatible performance to DROID-SLAM on average $t_{\mathrm{rel}}$, it outperforms all baselines in terms of $r_{\mathrm{rel}}$ by around $10\%$.
TABLE II: Performance comparison on the TartanAir v2 Hard Dataset. Noticeably, ORB-SLAM3 lost track of all sequences and is therefore not included. Only the hard subset is shown due to page limit, see Appendix. VI for the remaining results.
@@
      DROID-SLAM .0485 .1174 .0023 .0210 .0190 .0821 .0064 .0300 .0057 .0255 .1463 .2357 .0310 .0908 .0370 .0861
@@
      iSLAM-VO .4235 2.630 .3070 3.018 .3252 2.189 .3622 2.435 .2576 2.899 .2574 3.755 .2099 3.145 .3061 2.867
      TartanVO⋆† .1605 3.338 .2918 2.775 .2718 3.305 .2775 2.191 .2204 2.874 .1644 2.899 .2350 3.756 .2316 3.020
      TartanVO .1505 .4329 .0914 .7542 .0715 .3265 .0842 .4053 .0678 .6569 .0803 1.186 .0784 .9458 .0892 .6725
      DPVO⋆† .4984 .4937 .1738 .7112 .0539 .2539 .3847 2.703 .0481 .1869 .1891 1.430 .3365 2.943 .2406 1.246
Ours .0085 .1018 .0344 .1450 .0048 .0628 .0150 .0778 .0092 .1414 .0048 .0552 .0217 .4160 .0141 .1429
Monocular method.   Scale-aligned with ground truth.
TartanAir v2 Dataset    TartanAir v2 is challenging for visual SLAM. Our approach improves $61.9\%$ in $t_{\mathrm{rel}}$ compared to the nearest competitor. Notably, on trajectory H00, which simulates the lunar surface shown in Fig. 6, our model demonstrates a remarkable $82.4\%$ decrease in $t_{\mathrm{rel}}$ and achieves the lowest $r_{\mathrm{rel}}$ among all baseline methods.
TABLE III: Performance comparison of different methods on the KITTI Dataset. Only even numbered trajectory is shown due to page limit, see Appendix. VIII for remaining results.
@@
      ORB-SLAM 3 0.0252 0.0586 0.0438 0.0529 0.0274 0.0322 0.0228 0.0338 0.0271 0.0455 0.0166 0.0515 0.0258 0.0434
      DROID-SLAM 0.0198 0.0538 0.0250 0.0445 0.0255 0.0296 0.0199 0.0296 0.0275 0.0381 0.0309 0.0715 0.0900 0.0448
@@
      TartanVO⋆† 0.2066 0.1055 0.1626 0.1105 0.1152 0.0789 0.2234 0.0816 0.1857 0.0823 0.1745 0.0907 0.2207 0.0886
      TartanVO 0.0656 0.1026 0.0905 0.1197 0.1747 0.1158 0.0923 0.0968 0.0721 0.1063 0.0679 0.0969 0.1804 0.1147
      iSLAM-VO 0.0577 0.1052 0.0686 0.1101 0.1356 0.0787 0.0837 0.0812 0.0510 0.082 0.0449 0.0905 0.0878 0.0883
      DPVO⋆† 0.4542 0.0495 0.4209 0.0381 0.0348 0.0219 0.2393 0.0250 0.3051 0.0347 0.0661 0.0386 0.1951 0.0329
Ours 0.0192 0.0654 0.0223 0.0715 0.0206 0.0473 0.0187 0.0456 0.0254 0.0509 0.019 0.0569 0.0420 0.0645
Average over all trajectories (from 00 to 10) of KITTI Odom.   Monocular method.   Scale-aligned with ground truth.
KITTI Dataset    To further validate our robustness and consistency in outdoor, large-scale trajectory with the presence of dynamic objects, we evaluate our system on the KITTI [46] dataset. Our method, which relies solely on two-frame pose optimization without incorporating multi-frame bundle adjustment or loop closure, shows commendable performance, ranked behind the ORB-SLAM3, a full visual SLAM system, on $t_{\mathrm{rel}}$. Our system significantly outperforms other visual odometry approaches in $t_{\mathrm{rel}}$, demonstrating a 53.3% reduction in relative translation error. The performance observed in $r_{\mathrm{rel}}$ may be attributed to the lack of multi-frame optimization.
@@ IV-B Qualitative Analysis and Ablation Study
Figure 6: Top: Trajectories estimated by our model and baselines. We highlight the segments where dynamic objects interrupt the VO. Our method remains robust by implicitly filtering out unreliable features. Bottom left: We use our own platform with the ZED camera to collect data in the office. Our method demonstrates robustness against dynamic objects and visual occlusions in the images. Bottom right: samples of TartanAir v2 test dataset, which simulates the exotic lunar environment and factory with presence of dynamic objects.
In addition to quantitative evaluations, we conduct qualitative evaluations of our proposed system across multiple datasets including EuRoC, KITTI, and TartanAir v2, supplemented by manually collected data using a ZED stereo camera. Our model, even without multi-frame optimization, achieves top-tier performance and exhibits fewer glitches than baseline methods. As demonstrated in the Fig. 6, our method produces smoother trajectories and superior pose estimation precision. Fig. 6 a) presents that our model correctly identifies the region occluded by the mounted platform and dynamic objects in the scene and assigns a high uncertainty score to these regions.
Ablation Study   In Table. IV, we first perform the ablation study on each module of MAC-VO including (a. w/o CovKP & CovOPT) with random keypoint selector and identity covariance matrix. (b. w/o CovOpt) replace the metrics-aware covariance model with the identity covariance model. (c. w/o CovKP) replace the proposed keypoint selector with random keypoint selector.
To demonstrate the necessity of scale consistency and off-diagonal terms in our covariance model, we run the ablation study with different configurations: (I. DiagCov) remove off-diagonal terms in Eq. 3, (II. Scale-agnostic) normalize the covariance model by the average determinants of covariance matrices of each frame.
TABLE IV: Ablation study. Detailed results in Appendix. B-C.
@@
Figure 7: Results of various ablation setups on the trajectory H01 of TartanAir v2.
IV-C Runtime Analysis
The runtime analysis shown in Table. V uses the platform with AMD Ryzen 9 5950X CPU and NVIDIA 3090 Ti GPU. We also introduce a fast mode (MAC-VO Fast) that utilizes half-precision number in the network inference to enhance efficiency. This mode also speeds up the memory decoder network by reducing the number of iterative updates from 12 to 4. The fast mode performs at 10.5 fps (frames per second) with 70% of the performance of the original MAC-VO. More details are included in Appendix. C.
@@ V Conclusion & Discussion
This paper proposes MAC-VO, a learning-based stereo VO method that leverages the metrics-aware covariance model. Our model outperforms most visual odometry and even SLAM algorithms on challenging datasets. In our current work, the model focuses on the two-frame pose optimization. We believe our accuracy will be further benefit from bundle adjustment, multi-frame optimization, and loop closure. Additionally, we are interested in exploring our metrics-aware covariance model in multi-sensor fusion, such as with IMUs.
@@ A-A Depth Uncertainty From Disparity Uncertainty
A-B Depth Uncertainty with of Match Uncertainty
@@ A-C Projecting 2D Uncertainty to Spatial Covariance
Assume $\mathbf{u}_{i,t}$, $\mathbf{v}_{i,t}$, $\mathbf{d}_{i,t}$ are independent to each other, we have $\mathrm{Var}(\mathbf{u}_{i,t},\mathbf{d}_{i,t})=(\sigma_{u_{i,t}}^{2}+d_{i,t}^{2})(\sigma_{d_{i,t}}^{2}+u_{i,t}^{2})-u_{i,t}^{2}d_{i,t}^{2}$ [47]. Based on this expression of variance, it follows that
$\sigma_{x_{i,t}}^{2}$ $=\mathrm{Var}\left(\frac{\mathbf{u}_{i,t}\mathbf{d}_{i,t}}{f_{x}}-\frac{c_{x}\mathbf{d}_{i,t}}{f_{x}}\right)=\frac{\mathrm{Var}(\mathbf{u}_{i,t}\mathbf{d}_{i,t})}{f_{x}^{2}}+\frac{c_{x}^{2}\mathrm{Var}(\mathbf{d}_{i,t})}{f_{x}^{2}}$ (12)
$=\frac{(\sigma_{u_{i,t}}^{2}+d_{i,t}^{2})(\sigma_{d_{i,t}}^{2}+u_{i,t}^{2})-u_{i,t}^{2}d_{i,t}^{2}+c_{x}^{2}\sigma_{d_{i,t}}^{2}}{f_{x}^{2}}$
$\sigma_{y_{i,t}}^{2}$ $=\mathrm{Var}\left(\frac{\mathbf{v}_{i,t}\mathbf{d}_{i,t}}{f_{y}}-\frac{c_{y}\mathbf{d}_{i,t}}{f_{y}}\right)=\frac{\mathrm{Var}(\mathbf{v}_{i,t}\mathbf{d}_{i,t})}{f_{y}^{2}}+\frac{c_{y}^{2}\mathrm{Var}(\mathbf{d}_{i,t})}{f_{y}^{2}}$
$=\frac{(\sigma_{v_{i,t}}^{2}+d_{i,t}^{2})(\sigma_{d_{i,t}}^{2}+v_{i,t}^{2})-v_{i,t}^{2}d_{i,t}^{2}+c_{y}^{2}\sigma_{d_{i,t}}^{2}}{f_{y}^{2}}$
@@ B-A Remaining Results on EuRoC, TartanAirv2, and KITTI
TABLE VI: Performance comparison of different methods on the EuRoC Dataset. Only even-ordered trajectory is shown here, see Table. I for the remaining results.
@@
      ORB-SLAM 3 0.0036 0.0495 0.0061 0.0501 0.0049 0.0888 0.0137 0.2669 0.0090 0.1528
      DROID-SLAM 0.0012 0.0169 0.0031 0.0224 0.0024 0.0314 0.0036 0.0642 0.0017 0.0399
@@
      TartanVO⋆† 0.0172 0.0621 0.0213 0.0681 0.0124 0.0756 0.0263 0.1552 0.0171 0.1251
      TartanVO 0.0289 0.5037 0.0501 0.5400 0.0224 0.5322 0.0351 1.3127 0.0361 1.1607
      iSLAM-VO 0.0041 0.0620 0.0082 0.0682 0.0041 0.0756 0.0088 0.1554 0.0078 0.1252
      DPVO⋆† 0.0014 0.0212 0.0029 0.0264 0.0026 0.0405 0.0033 0.0662 0.0022 0.0493
Ours 0.0013 0.0199 0.0028 0.0273 0.0024 0.0304 0.0032 0.058 0.0018 0.0406
The estimated sequence is scale-aligned with ground truth.
Monocular method.
TABLE VII: Relative motion error on the TartanAir v2 Hard Dataset. Only even numbered trajectories in the Hard subset are shown here. For odd-numbered trajectories in the Hard subset, see Table. VIII. ORB-SLAM3 lost track of all trajectories and is therefore not shown in the table.
Trajectory H00 H02 H04 H06 H08 H10 H12
SLAM
      DROID-SLAM 0.009 0.033 0.029 0.080 0.014 0.026 0.006 0.025 0.019 0.082 0.004 0.060 0.146 0.236
VO
      iSLAM-VO 0.182 3.337 0.341 2.557 0.422 2.776 0.382 2.423 0.325 2.189 0.259 2.874 0.257 3.755
      TartanVO⋆† 0.097 0.459 0.315 2.629 0.247 3.084 0.234 3.017 0.272 3.305 0.245 2.435 0.164 2.899
      TartanVO 0.067 0.751 0.070 0.382 0.150 0.724 0.106 0.398 0.072 0.327 0.082 0.809 0.080 1.186
      DPVO⋆† 0.010 0.077 0.025 0.079 0.308 0.654 0.024 0.066 0.054 0.254 0.164 0.250 0.189 1.430
Ours 0.011 0.120 0.008 0.058 0.033 0.168 0.017 0.090 0.016 0.085 0.011 0.162 0.027 0.604
The estimated sequence is scale-aligned with ground truth.
Monocular method.
TABLE VIII: Performance comparison on the TartanAir v2 Easy Dataset.
@@
      ORB-SLAM3 .1019 2.349
      DROID-SLAM .0077 .0144 .0025 .0199 .0063 .0409 .0049 .0251 .0009 .0147 .0031 .0463 .0016 .0235 .0039 .0264
@@
      iSLAM-VO .0656 .2873 .0456 .3853 .0359 .2234 .0508 .2635 .0268 .3201 .0464 .6624 .0362 .4606 .0439 .3718
      TartanVO⋆† .0532 .3237 .0937 .4750 .1066 .6048 .0756 .2230 .1114 .4032 .0862 .3185 .1373 .6620 .0949 .4300
      TartanVO .0505 .1334 .0322 .2078 .0237 .1105 .0303 .1417 .0173 .1681 .0279 .3828 .0218 .2329 .0291 .1967
      DPVO⋆† .0113 .0187 .0047 .0249 .0099 .0475 .0603 .2064 .0044 .0177 .0511 .0665 .0189 .1543 .0229 .0766
Ours .0026 .0351 .0124 .1183 .0031 .0684 .0054 .0383 .0050 .0413 .0018 .0247 .0054 .1427 .0051 .0670
Monocular method.   Scale-aligned with ground truth.
TABLE IX: Performance comparison of different methods on the KITTI Dataset. Only odd-numbered trajectory is shown here, see Table. III for the remaining results.
@@
      ORB-SLAM 3 0.0416 0.0355 0.027 0.0425 0.0161 0.0416 0.0155 0.0385 0.0208 0.0444
      DROID-SLAM 0.7112 0.0406 0.0182 0.0385 0.0153 0.0353 0.0746 0.0734 0.0214 0.0378
@@
      TartanVO⋆† 0.6834 0.0895 0.1234 0.0682 0.1821 0.0761 0.2005 0.0847 0.1704 0.1069
      TartanVO 1.1408 0.2455 0.0477 0.0953 0.0637 0.0821 0.0700 0.0931 0.0990 0.1077
      iSLAM-VO 0.2978 0.0896 0.0507 0.0681 0.0504 0.0758 0.0593 0.0842 0.0660 0.1064
      DPVO⋆† 0.0942 0.0247 0.0302 0.0330 0.2221 0.0319 0.1064 0.0311 0.1723 0.0336
Ours 0.1670 0.1670 0.0504 0.0504 0.0466 0.0466 0.0507 0.0507 0.0567 0.0567
Average is calculated over all trajectories (from 00 to 10) of KITTI.
The estimated sequence is scale-aligned with ground truth.
Monocular method.

### v2

@@ IV-A Quantitative Analysis
TABLE I: Performance comparison of different methods on the EuRoC Dataset. Only odd-ordered trajectory is shown due to page limit, see Appendix B-A for the remaining results. Average is calculated over all trajectories of EuRoC.
@@
      [Uncaptioned image] ORB-SLAM 3 0.0035 0.0450 0.0058 0.0603 0.0059 0.0526 0.0096 0.1757 0.0064 0.1615 0.033 0.9497 0.0092 0.1866
      [Uncaptioned image] DROID-SLAM 0.0012 0.0159 0.0034 0.2656 0.0025 0.0193 0.0026 0.0417 0.0012 0.0289 0.0034 0.1033 0.0024 0.0590
      [Uncaptioned image] MASt3R-SLAM 0.0143 0.5493 0.0493 0.8890 0.0254 0.5808 0.0354 1.3178 0.0117 0.6818 0.0430 2.3306 0.0277 1.0472
@@
      [Uncaptioned image] TartanVO 0.0121 0.0560 0.0302 0.2791 0.0193 0.0604 0.0251 0.1244 0.0065 0.0920 0.0303 0.2986 0.0198 0.1270
      [Uncaptioned image] TartanVO 0.0277 0.5122 0.0514 0.6635 0.0464 0.4797 0.0394 1.0420 0.0195 0.4684 0.0473 1.9657 0.0368 0.8346
      [Uncaptioned image] iSLAM-VO 0.0042 0.0560 0.0076 0.2789 0.0070 0.0603 0.0066 0.1241 0.004 0.0920 0.0151 0.2984 0.0071 0.1269
      [Uncaptioned image] DPVO 0.0015 0.0207 0.0028 0.0273 0.0028 0.0243 0.0041 0.0496 0.0016 0.0342 0.0045 0.1205 0.0027 0.0437
      [Uncaptioned image] Ours 0.0014 0.0214 0.0023 0.0238 0.0025 0.0216 0.0029 0.0434 0.0012 0.0289 0.0049 0.1284 0.0024 0.0403
    [Uncaptioned image]   Monocular method.  [Uncaptioned image]   Stereo method.   Scale-aligned with ground truth.
EuRoC Dataset    We assessed our model on the EuRoC [48] dataset, as detailed in Table I, comparing it against baseline methods including visual odometry and state-of-the-art visual SLAM systems with loop-closure and global bundle adjustment. While our method exhibits compatible performance to DROID-SLAM on average $t_{\mathrm{rel}}$, it outperforms all baselines in terms of $r_{\mathrm{rel}}$ by around $10\%$.
TABLE II: Performance comparison on the TartanAir v2 Hard Dataset. Noticeably, ORB-SLAM3 and MASt3R-SLAM lost track on all sequences and are not included. See Appendix B-A for the Easy subset. Average is calculated over all trajectories of TartanAir v2 Hard.
@@
      [Uncaptioned image] DROID-SLAM .0485 .1174 .0023 .0210 .0190 .0821 .0064 .0300 .0057 .0255 .1463 .2357 .0310 .0908 .0370 .0861
@@
      [Uncaptioned image] TartanVO .1605 3.338 .2918 2.775 .2718 3.305 .2775 2.191 .2204 2.874 .1644 2.899 .2350 3.756 .2316 3.020
      [Uncaptioned image] TartanVO-stereo .1505 .4329 .0914 .7542 .0715 .3265 .0842 .4053 .0678 .6569 .0803 1.186 .0784 .9458 .0892 .6725
      [Uncaptioned image] iSLAM-VO .4235 2.630 .3070 3.018 .3252 2.189 .3622 2.435 .2576 2.899 .2574 3.755 .2099 3.145 .3061 2.867
      [Uncaptioned image] DPVO .4984 .4937 .1738 .7112 .0539 .2539 .3847 2.703 .0481 .1869 .1891 1.430 .3365 2.943 .2406 1.246
      [Uncaptioned image] Ours .0085 .1018 .0344 .1450 .0048 .0628 .0150 .0778 .0092 .1414 .0048 .0552 .0217 .4160 .0141 .1429
[Uncaptioned image]  Monocular method.  [Uncaptioned image]  Stereo method.   Scale-aligned with ground truth.
TABLE III: Performance comparison of different methods on the KITTI Dataset. Only even numbered trajectory is shown due to page limit, see Appendix B-A for remaining results. Average is calculated over all trajectories of KITTI Odometry.
@@
      [Uncaptioned image] ORB-SLAM 3 0.0252 0.0586 0.0438 0.0529 0.0274 0.0322 0.0228 0.0338 0.0271 0.0455 0.0166 0.0515 0.0258 0.0434
      [Uncaptioned image] DROID-SLAM 0.0198 0.0538 0.0250 0.0445 0.0255 0.0296 0.0199 0.0296 0.0275 0.0381 0.0309 0.0715 0.0900 0.0448
      [Uncaptioned image] MASt3R-SLAM 0.8988 0.7154 1.0104 0.5959 1.2266 0.1124 1.2599 0.1289 0.7760 0.6075 0.6889 0.4709 0.9538 0.4202
@@
      [Uncaptioned image] TartanVO 0.2066 0.1055 0.1626 0.1105 0.1152 0.0789 0.2234 0.0816 0.1857 0.0823 0.1745 0.0907 0.2207 0.0886
      [Uncaptioned image] TartanVO 0.0656 0.1026 0.0905 0.1197 0.1747 0.1158 0.0923 0.0968 0.0721 0.1063 0.0679 0.0969 0.1804 0.1147
      [Uncaptioned image] iSLAM-VO 0.0577 0.1052 0.0686 0.1101 0.1356 0.0787 0.0837 0.0812 0.0510 0.082 0.0449 0.0905 0.0878 0.0883
      [Uncaptioned image] DPVO 0.4542 0.0495 0.4209 0.0381 0.0348 0.0219 0.2393 0.0250 0.3051 0.0347 0.0661 0.0386 0.1951 0.0329
      [Uncaptioned image] Ours 0.0192 0.0654 0.0223 0.0715 0.0206 0.0473 0.0187 0.0456 0.0254 0.0509 0.019 0.0569 0.0420 0.0645
[Uncaptioned image]  Monocular method.  [Uncaptioned image]  Stereo method.   Scale-aligned with ground truth.
TartanAir v2 Dataset    TartanAir v2 presents significant challenges for visual odometry due to its rigorous motion and extreme illumination changes. The traditional method, ORB-SLAM, fails on most trajectories. Our approach improves $61.9\%$ in $t_{\mathrm{rel}}$ compared DroidSLAM. Notably, on trajectory H00, which simulates the lunar surface shown in Fig. 6, our model demonstrates a remarkable $82.4\%$ decrease in $t_{\mathrm{rel}}$ and achieves the lowest $r_{\mathrm{rel}}$ among all baseline methods.
KITTI Dataset   To further validate our robustness and consistency in outdoor, large-scale trajectory with dynamic objects, we evaluate our system on the KITTI [49] dataset. Our method, which relies solely on two-frame pose optimization, shows commendable performance, ranked behind the ORB-SLAM3, a full visual SLAM system, on $t_{\mathrm{rel}}$. Our system significantly outperforms other visual odometry approaches in $t_{\mathrm{rel}}$, demonstrating a 53.3% reduction in relative translation error. The degraded performance of $r_{\mathrm{rel}}$ may be attributed to the lack of multi-frame optimization.
@@ IV-B Qualitative Analysis and Ablation Study
Figure 6: Top: Trajectories estimated by our model and baselines. We highlight the segments where dynamic objects interrupt the VO. Bottom left: We collect data in the office using our own payload with the ZED-X camera. The data contains multiple dynamic objects and visual occlusions. Bottom right: samples of TartanAir v2 test dataset, which simulates the exotic lunar environment.
In addition to quantitative evaluations, we conduct qualitative evaluations of our proposed system across multiple datasets including EuRoC, KITTI, and TartanAir v2, supplemented by manually collected data using a ZED-X stereo camera. MASt3R-SLAM struggles on many trajectories of KITTI and TartanAir, primarily due to incorrect loop closure detections from visual ambiguities. Our model, even without multi-frame optimization, achieves top-tier performance and exhibits fewer glitches than baseline methods. As demonstrated in the Fig. 6, our method produces smoother trajectories and superior pose estimation precision. Fig. 6 a) presents that our model correctly identifies the region occluded by the mounted platform and dynamic objects in the scene and assigns a high uncertainty score to these regions.
TABLE IV: Ablation study. Detailed results in Appendix B-C.
@@
Figure 7: Results of various ablation setups on the H01 of TartanAirv2.
Ablation Study   In Table IV, we first perform the ablation study on each module of MAC-VO including (a. w/o CovKP & CovOPT) with random keypoint selector and identity covariance matrix. (b. w/o CovOpt) replace the metrics-aware covariance model with the identity covariance model. (c. w/o CovKP) replace the proposed keypoint selector with a random keypoint selector. To demonstrate the necessity of scale consistency and off-diagonal terms in our covariance model, we run the ablation study with different configurations: (I. DiagCov) remove off-diagonal terms in Eq. 3, (II. Scale-agnostic) normalize the covariance model by the average determinants of covariance matrices of each frame.
@@
IV-C Runtime Analysis
The runtime analysis shown in Table V uses the platform with AMD Ryzen 9 5950X CPU and NVIDIA 3090 Ti GPU. We also introduce a fast mode (MAC-VO Fast) that utilizes half-precision number in the network inference to enhance efficiency. This mode also speeds up the memory decoder network by reducing the number of iterative updates from 12 to 4. The fast mode performs at 10.5 fps (frames per second) with 70% of the performance of the original MAC-VO. More details are included in Appendix C.
@@ V Conclusion & Discussion
This paper proposes MAC-VO, a learning-based stereo VO method that outperforms most visual odometry and even SLAM algorithms on challenging datasets. In our current work, the model focuses on the two-frame pose optimization. We believe our accuracy will be further benefit from bundle adjustment, multi-frame optimization, and loop closure. Additionally, we plan to apply our metrics-aware covariance model in multi-sensor fusion, such as with IMUs.
@@ A-A Depth Uncertainty From Disparity Uncertainty
A-B Uncertainty Correction After Keypoint Matching
@@ A-C Projecting 2D Uncertainty to Spatial Covariance
Assume $\mathbf{u}_{i,t}$, $\mathbf{v}_{i,t}$, $\mathbf{d}_{i,t}$ are independent to each other, we have $\mathrm{Var}(\mathbf{u}_{i,t}\mathbf{d}_{i,t})=(\sigma_{u_{i,t}}^{2}+d_{i,t}^{2})(\sigma_{d_{i,t}}^{2}+u_{i,t}^{2})-u_{i,t}^{2}d_{i,t}^{2}$ [50]. Based on this expression of variance, it follows that
$\sigma_{x_{i,t}}^{2}$ $=\mathrm{Var}\left(\frac{(\mathbf{u}_{i,t}-c_{x})\mathbf{d}_{i,t}}{f_{x}}\right)=\frac{\mathrm{Var}((\mathbf{u}_{i,t}-c_{x})\mathbf{d}_{i,t})}{f_{x}^{2}}$ (12)
$=\frac{\sigma_{u_{i,t}}^{2}\sigma_{d_{i,t}}^{2}+\sigma_{u_{i,t}}^{2}\mathbf{d}_{i,t}^{2}+(\mathbf{u}_{i,t}-c_{x})^{2}\sigma_{d_{i,t}}^{2}}{f_{x}^{2}}$
$\sigma_{y_{i,t}}^{2}$ $=\mathrm{Var}\left(\frac{(\mathbf{v}_{i,t}-c_{y})\mathbf{d}_{i,t}}{f_{y}}\right)=\frac{\mathrm{Var}((\mathbf{v}_{i,t}-c_{y})\mathbf{d}_{i,t})}{f_{y}^{2}}$
$=\frac{\sigma_{v_{i,t}}^{2}\sigma_{d_{i,t}}^{2}+\sigma_{v_{i,t}}^{2}\mathbf{d}_{i,t}^{2}+(\mathbf{v}_{i,t}-c_{y})^{2}\sigma_{d_{i,t}}^{2}}{f_{y}^{2}}$
@@ B-A Remaining Results on EuRoC, TartanAirv2, and KITTI
See Table VI, Table VII, and Table VIII for data.
TABLE VI: Performance comparison of different methods on the EuRoC Dataset. Only even-ordered trajectory is shown here, see Table I for the remaining results.
@@
      [Uncaptioned image] ORB-SLAM 3 0.0036 0.0495 0.0061 0.0501 0.0049 0.0888 0.0137 0.2669 0.0090 0.1528
      [Uncaptioned image] DROID-SLAM 0.0012 0.0169 0.0031 0.0224 0.0024 0.0314 0.0036 0.0642 0.0017 0.0399
      [Uncaptioned image] MASt3R-SLAM 0.0138 0.5493 0.0274 0.8890 0.0130 0.5808 0.0383 1.3178 0.0334 0.6818
@@
      [Uncaptioned image] TartanVO 0.0172 0.0621 0.0213 0.0681 0.0124 0.0756 0.0263 0.1552 0.0171 0.1251
      [Uncaptioned image] TartanVO 0.0289 0.5037 0.0501 0.5400 0.0224 0.5322 0.0351 1.3127 0.0361 1.1607
      [Uncaptioned image] iSLAM-VO 0.0041 0.0620 0.0082 0.0682 0.0041 0.0756 0.0088 0.1554 0.0078 0.1252
      [Uncaptioned image] DPVO 0.0014 0.0212 0.0029 0.0264 0.0026 0.0405 0.0033 0.0662 0.0022 0.0493
      [Uncaptioned image] Ours 0.0013 0.0199 0.0028 0.0273 0.0024 0.0304 0.0032 0.058 0.0018 0.0406
[Uncaptioned image]  Monocular method.  [Uncaptioned image]  Stereo method.   Scale-aligned with ground truth.
TABLE VII: Performance comparison on the TartanAir v2 Easy Dataset.
@@
      [Uncaptioned image] ORB-SLAM3 - - .1019 2.349 - - - - - - - - - - .1019 2.349
      [Uncaptioned image] MASt3R-SLAM .3268 2.892 .3124 2.921 - - - - .1872 2.948 - - .2018 3.033 .2570 .2948
      [Uncaptioned image] DROID-SLAM .0077 .0144 .0025 .0199 .0063 .0409 .0049 .0251 .0009 .0147 .0031 .0463 .0016 .0235 .0039 .0264
@@
      [Uncaptioned image] TartanVO .0532 .3237 .0937 .4750 .1066 .6048 .0756 .2230 .1114 .4032 .0862 .3185 .1373 .6620 .0949 .4300
      [Uncaptioned image] TartanVO .0505 .1334 .0322 .2078 .0237 .1105 .0303 .1417 .0173 .1681 .0279 .3828 .0218 .2329 .0291 .1967
      [Uncaptioned image] iSLAM-VO .0656 .2873 .0456 .3853 .0359 .2234 .0508 .2635 .0268 .3201 .0464 .6624 .0362 .4606 .0439 .3718
      [Uncaptioned image] DPVO .0113 .0187 .0047 .0249 .0099 .0475 .0603 .2064 .0044 .0177 .0511 .0665 .0189 .1543 .0229 .0766
      [Uncaptioned image] Ours .0026 .0351 .0124 .1183 .0031 .0684 .0054 .0383 .0050 .0413 .0018 .0247 .0054 .1427 .0051 .0670
[Uncaptioned image]  Monocular method.  [Uncaptioned image]  Stereo method.   Scale-aligned with ground truth.
TABLE VIII: Performance comparison of different methods on the KITTI Dataset. Only odd-numbered trajectory is shown here, see Table III for the remaining results.
@@
      [Uncaptioned image] ORB-SLAM 3 0.0416 0.0355 0.027 0.0425 0.0161 0.0416 0.0155 0.0385 0.0208 0.0444
      [Uncaptioned image] DROID-SLAM 0.7112 0.0406 0.0182 0.0385 0.0153 0.0353 0.0746 0.0734 0.0214 0.0378
      [Uncaptioned image] MASt3R-SLAM 1.9316 0.3442 0.5313 0.2917 - - 0.3735 0.3913 0.8414 0.5479
@@
      [Uncaptioned image] TartanVO 0.6834 0.0895 0.1234 0.0682 0.1821 0.0761 0.2005 0.0847 0.1704 0.1069
      [Uncaptioned image] TartanVO 1.1408 0.2455 0.0477 0.0953 0.0637 0.0821 0.0700 0.0931 0.0990 0.1077
      [Uncaptioned image] iSLAM-VO 0.2978 0.0896 0.0507 0.0681 0.0504 0.0758 0.0593 0.0842 0.0660 0.1064
      [Uncaptioned image] DPVO 0.0942 0.0247 0.0302 0.0330 0.2221 0.0319 0.1064 0.0311 0.1723 0.0336
      [Uncaptioned image] Ours 0.2663 0.1670 0.0187 0.0504 0.0142 0.0466 0.0194 0.0507 0.0182 0.0567
[Uncaptioned image]  Monocular method.  [Uncaptioned image]  Stereo method.   Scale-aligned with ground truth.